OpenStudy (pawanyadav):

Let f(x)=| |e^x-1| -1| then f(x) is non differentlable for x belongs to:

OpenStudy (tkhunny):

Show your work, please. What have you done, so far?

OpenStudy (loser66):

@tkhunny I would like to know how to solve it algebraically also. Please, guide me. It doesn't matter if the Asker doesn't do his work. There are plenty of other students like me take advantage on other's stuff.

OpenStudy (tkhunny):

First, we must establish the meaning of the notation. Is it \(e^{x-1}\;or\;e^{x}-1\)? The latter might be more interesting.

OpenStudy (loser66):

oh, i put |(e^x) -1| -1| on graphing tool to see the point the function is nondifferentiable but not know how to find it :)

OpenStudy (tkhunny):

Where is \(e^{x}-1 \ge 0\)? Where is \(e^{x}-1 < 0\)?

OpenStudy (loser66):

when x > 0 when x <0

OpenStudy (loser66):

but |e^x -1| > 0 for all x but at x =0 we have a cups there.

OpenStudy (pawanyadav):

y=|e^x-1|-1 try to plot its graph

OpenStudy (pawanyadav):

You will get upward parabola from (0,-1) And (ln2,0) in 1 st and 4 the quadrant This part of graph is sufficient for the question So, f(x) is non-differentiable for x belongs to {0,ln2} ,..that's what the question asking.

OpenStudy (pawanyadav):

@Loser66

OpenStudy (pawanyadav):

@tkhunny this is the work

OpenStudy (loser66):

https://www.desmos.com/calculator/r6uvvc16xs

OpenStudy (loser66):

only (0,1) and (0.693, 0) are two point the graph is nondiffentiable.

OpenStudy (pawanyadav):

Yes and this is also the answer.

OpenStudy (loser66):

but I don't know how to solve it algebraically. We need it to back the logic up. ok, let try one. We need find the root of the function ||e^x -1 |-1| =0 when

OpenStudy (pawanyadav):

Dude, solving these problems graphically is the best method.

OpenStudy (loser66):

I was not taught that!! to my courses, we have to prove it algebraically, graph is just to illustrate what we do.

OpenStudy (pawanyadav):

But it saves your time OK lets do iit

OpenStudy (pawanyadav):

How will you starstart

OpenStudy (pawanyadav):

|e^x-1|-1=0

OpenStudy (pawanyadav):

So |e^x-1 |= 1 Now two case e^x-1=1. OR. e^x-1=-1 e^x=2. Or. e^x=0

OpenStudy (pawanyadav):

I think you can solve from there

OpenStudy (pawanyadav):

Do you need more help

OpenStudy (tkhunny):

You did not answer my questions. Where is \(e^{x} − 1 \ge 0\) ? Where is \(e^{x} − 1 < 0\) ? No need to draw pictures. If you ALWAYS have to draw pictures, what will you do in 4 dimensions?

OpenStudy (pawanyadav):

Every question have different ways to do....I do them in the way they suits me.

OpenStudy (pawanyadav):

e^x-1>=0 when x>=0 And e^x-1<0 when x<0

OpenStudy (pawanyadav):

@tkhunny

OpenStudy (tkhunny):

We have \(f(x) = ||e^{x}-1|-1|\) When x >= 0, we have \(f_{1}(x) = |(e^{x}-1)-1| = |e^{x} - 2|\) When x < 0, we have \(f_{2}(x) = |(1 - e^{x})-1| = |e^{x}| = e^{x}\) We now KNOW something is funny at x = 0. This is a good place to check for differentiability. Then we ask again, when is \(e^{x} - 2 \ge 0\) and when is \(e^{x} - 2 < 0\)? We will find more suspicion associated with this answer.

OpenStudy (anonymous):

x in reals only?

OpenStudy (pawanyadav):

e^x-2>=0 e^x>=2 By taking ln we get x>=ln2

OpenStudy (pawanyadav):

And e^x<2 When x<ln2 @tkhunny

OpenStudy (pawanyadav):

Means something is also funny at x=ln2 So we should also check differentiability here So we get two points x=0 and x=ln2

OpenStudy (pawanyadav):

@tkhunny

OpenStudy (tkhunny):

Check it out! There is no reason to believe it is not differentiable elsewhere. Feel free to write it out: \(f(x) = \begin{cases} e^{x} & x < 0 \\ 2 - e^{x} & 0 \le x < \ln(2) \\ e^{x} - 2 & \ln(2) \le x \ \end{cases} \)

OpenStudy (pawanyadav):

This method is time consuming than graph isn't it?

OpenStudy (pawanyadav):

I can't do the question fastly with this in exam

OpenStudy (pawanyadav):

Can you taught me how to find no. of solution of a function?

OpenStudy (tkhunny):

That would depend on the speed of your thought. Think fundamentals FIRST - THEN worry about speed. Personally, in my college career, I never worried about speed. I concerned myself ONLY with fundamentals and I still managed to finish exams faster than most and with higher scores than most. If you are going to plan for speed, you must be very careful not to become sloppy.

OpenStudy (tkhunny):

You cannot find "ln(2)" on your graph. Will 0.693147 constitute a sufficient response? Maybe.

OpenStudy (pawanyadav):

Sin^-1 x-|x-a|=0 ,is having at least one real root,then possible values of a can be?

OpenStudy (pawanyadav):

But for entering a college, I have to crack an exam in which I got only 3 minutes for such questions.

OpenStudy (tkhunny):

Generally, one question per post. Always show YOUR work first - right up front.

OpenStudy (pawanyadav):

OK wai

OpenStudy (tkhunny):

Easier to find if you are looking. Also, some folks care about medals. You can give only one, no matter how long the thread.

OpenStudy (pawanyadav):

I draw graph sin^-1 x.. Then y=a-x passed through point (1,π/2)

OpenStudy (pawanyadav):

Can you have some good approache for it the way I tried is not much working

OpenStudy (pawanyadav):

@tkhunny

OpenStudy (pawanyadav):

What if I do like this Sin ^-1(x)=|x-a| -π/2<(x-a)>π/2 or -π/2<(a-x)>π/2

OpenStudy (tkhunny):

Inverse sine is limited to [-1,+1] We're talking about staying withing '1' of 'a'.

OpenStudy (pawanyadav):

What will I do next